# Deformations of the Fano scheme of a cubic

### v2

@@ Introduction
Let $\mathbf{P}$ be the complex projective space of dimension $d+1$, and $\mathrm{X}\subset\mathbf{P}$ a cubic with a finite number of singularities. For $d\geqslant 3$, it is well-known that the geometry of $\mathrm{X}$ is largely determined by the Hilbert scheme $\mathrm{F}=\mathrm{F}(\mathrm{X})$ of lines on $\mathrm{X}$, which is traditionally called the Fano scheme of $\mathrm{X}$. A great deal is known about $\mathrm{F}$ for $d=3$ or $d=4$ [2, 9, 11, 14, 29], and so our focus is on the $d\geqslant 5$ case, which has received much less attention. Altman and Kleiman [1] show that $\mathrm{F}$ is an irreducible normal local complete intersection of dimension $42d-$ 4, and it is known that $\mathrm{X}$ can be recovered from $\mathrm{F}$ [8].
@@
[…]
@@
${\mathrm{H}^{0}(\mathcal{N}_{\mathrm{X}/\mathbf{G}})}$${\operatorname{Ext}^{1}(\Omega^{1}_{\mathrm{X}},\mathcal{O}_{\mathrm{X}})}$${\mathrm{H}^{0}(\mathcal{N}_{\mathrm{F}/\mathbf{G}})}$${\operatorname{Ext}^{1}(\Omega^{1}_{\mathrm{F}},\mathcal{O}_{\mathrm{F}})}$$\scriptstyle{d\eta_{\mathcal{H}}}$$\scriptstyle{d\eta}$
of differentials. Relying on Borel-Bott-Weil computations and hypercohomology spectral sequences associated to the Koszul resolution of $\mathcal{O}_{\mathrm{F}}$, we show that $\mathrm{H}^{0}(\mathcal{N}_{\mathrm{F}/\mathbf{G}})\rightarrow\operatorname{Ext}^{1}(\Omega^{1}_{\mathrm{F}},\mathcal{O}_{\mathrm{F}})$ and $d\eta_{\mathcal{H}}$ are surjective; we then observe that
@@
using a result of Charles [8] which relates the automorphism groups of $\mathrm{F}$ to the one of $\mathrm{X}$. The condition $\mathrm{H}^{0}(\Theta_{\mathrm{X}})=0$, which holds for example for Lefschetz cubics, then guarantees that both $\mathcal{D}_{\mathrm{X}}$ and $\mathcal{D}_{\mathrm{F}}$ are pro-representable. Our proof shows that, without assuming $\mathrm{H}^{0}(\Theta_{\mathrm{X}})=0$, $\eta_{\mathcal{H}}$ is an isomorphism and $\eta$ is surjective.
We should discuss the relation of our functorial approach to the work of Borcea [5, 6] and Wehler [29]. Writing $\mathrm{X}=\mathrm{Z}(f)$ for $f\in\mathrm{H}^{0}(\mathcal{O}_{\mathbf{P}}(3))$, Borcea [6] considers the deformation of $\mathrm{F}$ given by varying $f$ in $\mathrm{H}^{0}(\mathcal{O}_{\mathbf{P}}(3))$. He checks the conditions
@@
which guarantee the completeness of the deformation [6, 29], for $d\geqslant 6$. In contrast to his and other papers [1, 10] using similar methods, we explicitly compute the decomposition of the sheaves $\Lambda^{n}\mathrm{S}^{3}\mathcal{S}$ (which occur in the Koszul resolution of $\mathcal{J}_{\mathrm{F}/\mathbf{G}}$) into Schur powers. This allows us to check the conditions (1), which play an important role in our proof, for all $d\geqslant 5$, thus extending Borcea’s result to $d=5$. In Theorem 3.1, we use the decomposition of $\Lambda^{n}\mathrm{S}^{3}\mathcal{S}$ to express the Hilbert polynomial $\chi(\mathcal{O}_{\mathrm{F}}(n))$ of $\mathrm{F}$ in terms of the Pochhammer symbol; this generalises previous results of Altman and Kleiman [1] and Libgober [19].
@@
of locally free sheaves. The Borel-Bott-Weil theorem, which we will use frequently in this paper, computes the cohomology of sheaves of the form
@@
Theorem 2.1 (Borel-Bott-Weil [4]).
@@
We will in particular rely on the following standard applications, where we use tacitly the canonical isomorphism $\mathrm{S}^{\vee}=\mathcal{S}\otimes\det(\mathcal{Q})$.
@@
and write $q_{\mathrm{S}}$ and $p_{\mathrm{S}}$ for the projections of $\mathbf{P}_{\mathrm{S}}\times_{\mathrm{S}}\mathrm{F}(\mathbf{P}_{\mathrm{S}}/\mathrm{S})$ to $\mathbf{P}_{\mathrm{S}}$ and $\mathrm{F}(\mathbf{P}_{\mathrm{S}}/\mathrm{S})$, respectively. By Theorem 2.17 of [1], the closed subscheme $\mathrm{F}(\mathrm{X}/\mathrm{S})\subset\mathrm{F}(\mathbf{P}_{\mathrm{S}}/\mathrm{S})$ is the zero scheme of the canonical morphism
@@
of sheaves on $\mathbf{P}_{\mathrm{S}}\times_{\mathrm{S}}\mathrm{F}(\mathbf{P}_{\mathrm{S}}/\mathrm{S})$. Of course, $\mathrm{F}(\mathbf{P}_{\mathrm{S}}/\mathrm{S})=\mathbf{G}\times\mathrm{S}$, where we view $\mathbf{G}$ as $\mathrm{F}(\mathbf{P})$; writing $\pi:\mathbf{G}\times\mathrm{S}\rightarrow\mathrm{S}$ for the projection, we have
@@
If $\mathrm{X}=\mathrm{Z}(f)$ for $f\in\mathrm{H}^{0}(\mathcal{O}_{\mathbf{P}_{\mathrm{S}}}(3))$, then applying $p_{\mathrm{S}}$ to (4) induces a section $\sigma_{f}$ of
@@
such that, invoking Proposition 2.3 of [1],
@@
If $\mathrm{S}=\operatorname{Spec}(\Lambda)$ is affine, we use the abbreviation $\mathrm{F}(\mathrm{X}/\Lambda)=\mathrm{F}(\mathrm{X}/\operatorname{Spec}(\Lambda))$. We first consider $\mathrm{F}(\mathrm{X})=\mathrm{F}(\mathrm{X}/\mathbf{C})$ for a cubic $\mathrm{X}=\mathrm{Z}(f)$, $f\in\mathrm{H}^{0}(\mathcal{O}_{\mathbf{P}}(3))$. This scheme is particularly well-behaved when the singular locus of $\mathrm{X}$ is finite:
Theorem 2.2 (Altman-Kleiman [1]).
Let $\mathrm{X}$ be a cubic with finitely many singularities. The Hilbert scheme $\mathrm{F}=\mathrm{F}(\mathrm{X})$ is of pure dimension $42d-$ 4; moreover, $\mathrm{F}$ is reduced for $d\geqslant 4$.
@@
Here the sum is taken over all partitions $\lambda$ of $nm$ with at most $n$ parts, and the numbers $a^{\lambda}_{n,m}$ can be expressed in terms of generalized Kostka numbers [20]. For small $n$ and $m$, these coefficients are relatively easy to compute; we find
@@
By applying Lemma 2.1, $\mathcal{S}^{\vee}=\mathcal{S}\otimes\det(\mathcal{Q})$, and the Pieri rule, we obtain
@@
For certain classes of complete intersections (including cubics of dimension $d\geqslant 6$), the latter result was obtained by Borcea [5]; our approach is similar to his, but Borcea does not explicitly compute the plethysms of Lemma 2.1 — by employing weight considerations, he instead proves a vanishing theorem (which, by Proposition 2.1, does not hold for $d=5$).
@@
We recall now some well-known general facts about functors of Artin rings [22, 24], and explain our notation; for us an Artin ring is a local $\mathbf{C}$-algebra which is finite over $\mathbf{C}$. For a functor of Artin rings $\mathcal{F}$, we denote by $t_{\mathcal{F}}=\mathcal{F}(\mathbf{C}[\varepsilon])$ the tangent space of $\mathcal{F}$, and if $\varphi:\mathcal{F}\rightarrow\mathcal{G}$ is a functorial morphism, we refer to
@@
as the differential of $\varphi$. For future reference, we recall the following general result [24]:
@@
The local moduli functor $\mathcal{D}_{\mathrm{S}}$ of a projective scheme $\mathrm{S}$ takes an Artin ring $\Lambda$ to the set $\mathcal{D}_{\mathrm{S}}(\Lambda)$ of isomorphism classes of deformations of $\mathrm{S}$ over $\Lambda$. A basic result is:
@@
(iii) If $\mathrm{S}$ is reduced, then $t_{\mathcal{D}_{\mathrm{S}}}=\operatorname{Ext}^{1}(\Omega^{1}_{\mathrm{S}},\mathcal{O}_{\mathrm{S}})$; if $\mathrm{S}$ also a local complete intersection, then $\operatorname{Ext}^{2}(\Omega^{1}_{\mathrm{S}},\mathcal{O}_{\mathrm{S}})$ is an obstruction space for $\mathcal{D}_{\mathrm{S}}$.
Here (i) is a theorem of Schlessinger [22], while (ii) goes back to Kodaira and Spencer [17] (see also [12]). For a closed subscheme $\mathrm{Z}$ of $\mathrm{S}$, the Hilbert functor of $\mathrm{S}$ induces a functor of Artin rings (local Hilbert functor) $\mathcal{H}_{\mathrm{Z}/\mathrm{S}}$, which takes an Artin ring $\Lambda$ to the set $\mathcal{H}_{\mathrm{Z}/\mathrm{S}}(\Lambda)$ of deformations of $\mathrm{Z}$ in $\mathrm{S}$ over $\Lambda$. By the existence of the Hilbert scheme of $\mathrm{S}$, $\mathcal{H}_{\mathrm{Z}/\mathrm{S}}$ is pro-representable and $t_{\mathcal{H}_{\mathrm{Z}/\mathrm{S}}}=\mathrm{H}^{0}(\mathcal{N}_{\mathrm{Z}/\mathrm{S}})$. It is related to $\mathcal{D}_{\mathrm{Z}}$ by a forgetful morphism
@@
We now assume that $\mathrm{S}$ is smooth, and $\mathcal{F}$ a locally free sheaf on $\mathrm{S}$.
@@
Finally, we consider the closed subscheme $\mathcal{Z}$ of $\mathrm{H}^{0}(\mathcal{F})\times\mathrm{S}$, which parametrises pairs $(\sigma,s)$ with $\sigma(s)=0$. The fibre of the projection
@@
over $\sigma\in\mathrm{H}^{0}(\mathcal{F})$ is the zero scheme $\mathrm{Z}=\mathrm{Z}(\sigma)$. If $\sigma$ is regular, then $\pi$ is flat in a neighbourhood of $\sigma$, thus inducing a deformation of $\mathrm{Z}$. The following result gives a criterion for the completeness of the latter deformation.
Lemma 2.4.
@@
Using Schubert calculus, Altman and Kleiman [1] prove that the Plücker degree of $\mathrm{F}$ is
@@
In the special case $d=3$, this is a theorem of Fano [11]. It is thus a natural question to determine, more generally, the Hilbert polynomial
$\chi(\mathcal{O}_{\mathrm{F}}(n))=\sum_{k=0}^{2d-4}\frac{n^{k}}{k!}\int_{\mathrm{F}}c_{1}(\mathcal{O}_{\mathrm{F}}(1))^{k}\cap\mathrm{Td}(\mathrm{F})$
of $\mathrm{F}$. Altman and Kleiman (and, independently, Libgober [19]) show that for $d=3$, we have
@@
In this section, we use LABEL:plethysm to express the Hilbert polynomial $\chi(\mathcal{O}_{\mathrm{F}}(n))$, for any dimension $d$, in terms of the Pochhammer symbol.
@@
Using Lemma 2.1, it suffices to describe the Hilbert polynomial of $\Sigma^{\mu_{1},\mu_{2}}\mathcal{S}$ for any $\mu_{1}\geqslant\mu_{2}$. We now establish the equality
@@
$\chi(\Sigma^{n^{d}}\mathcal{Q}\otimes\Sigma^{\mu_{1},\mu_{2}}\mathcal{S})=\dim\Sigma^{n^{d},\mu_{1},\mu_{2}}\mathrm{V}.$
@@
by the Weyl dimension formula. In particular, combining (7) with Lemma 2.1, we obtain
@@
A result of Schlömlich [23] explicitly describes the coefficients of
@@
[${d\brack k}=(-1)^{d-k}\sum_{m=0}^{d-k}\binom{d-1+m}{k-1}\binom{2d-k}{d+m}\sum_{n=0}^{m}\frac{(-1)^{n}n^{d-k+m}}{n!(m-n)!}.$
Applying this to Theorem 3.1, we can express the coefficients of $\chi(\mathcal{O}_{\mathrm{F}}(n))$ as sums of binomial coefficients; this shows in particular that our expression for $\chi(\mathcal{O}_{\mathrm{F}}(n))$ is a polynomial of degree $42d-$ 4, as it should.
@@
Theorem 3.1 and Kodaira vanishing
@@
allow one to compute the dimension of the space $\mathrm{H}^{0}(\mathcal{J}_{\mathrm{F}/\mathbf{G}}(n))$ of global sections of $\mathcal{O}_{\mathbf{G}}(n)$ vanishing on $\mathrm{F}$. Indeed, for $d\geqslant 4$, Debarre and Manivel [10] prove that $\mathrm{H}^{1}(\mathcal{J}_{\mathrm{F}/\mathbf{G}}(n))=0$ for $n\geqslant 0$. There is thus an exact sequence of the form.
$0\rightarrow\mathrm{H}^{0}(\mathcal{J}_{\mathrm{F}/\mathbf{G}}(n))\rightarrow\mathrm{H}^{0}(\mathcal{O}_{\mathbf{G}}(n))\rightarrow\mathrm{H}^{0}(\mathcal{O}_{\mathrm{F}}(n))\rightarrow 0$
@@
Writing $\mathrm{X}_{d}$ to indicate the dimension $d$ of the cubic $\mathrm{X}$, combining Theorem 3.1 with Schlömlich’s formula gives
@@
In fact, any deformation $\mathfrak{X}\subset\mathbf{P}_{\Lambda}$ of $\mathrm{X}$ in $\mathbf{P}$ over an Artin ring $\Lambda$ is a cubic: there exists a section $f_{\Lambda}$ of $\mathcal{O}_{\mathbf{P}_{\Lambda}}(3)$ extending $f$, such that $\mathfrak{X}=\mathrm{Z}(f_{\Lambda})$ [17, 30].
@@
As the vanishing of $\mathrm{H}^{0}(\Theta_{\mathrm{X}})$ guarantees the pro-representability of $\mathcal{D}_{\mathrm{X}}$, we are led to study $\mathrm{H}^{0}(\Theta_{\mathrm{X}})$. It is well-known that $\mathrm{H}^{0}(\Theta_{\mathrm{X}})=0$ when $\mathrm{X}$ is smooth [15]. We observe here that this holds for the simplest class of singular cubics: Lefschetz cubics in the sense of [9], i.e. those with at most one node. Since $\mathrm{H}^{0}(\Theta_{\mathrm{X}})$ is the kernel of the derivative
@@
${\mathrm{H}^{0}(\Theta_{\mathrm{P}}|_{\mathrm{X}})}$${\mathrm{H}^{0}(\mathcal{O}_{\mathrm{X}}(3))}$${\mathrm{H}^{0}(\Theta_{\mathrm{P}})}$${\mathrm{H}^{0}(\mathcal{O}_{\mathrm{P}}(3))/(f)}$$\scriptstyle{df}$
[…]
is given by $df(\sum L_{i}\partial_{i})=\sum L_{i}\partial_{i}f\mod(f)$. We can thus view $\mathrm{H}^{0}(\Theta_{\mathrm{X}})$ as the subspace of $\mathrm{H}^{0}(\Theta_{\mathrm{P}})$ consisting of all $\sum L_{i}\partial_{i}$ such that

### v3

@@ Introduction
Let $\mathbf{P}$ be the complex projective space of dimension $d+1$, and $\mathrm{X}\subset\mathbf{P}$ a cubic with a finite number of singularities. For $d\geqslant 3$, it is well-known that the geometry of $\mathrm{X}$ is largely determined by the Hilbert scheme $\mathrm{F}=\mathrm{F}(\mathrm{X})$ of lines on $\mathrm{X}$, which is traditionally called the Fano scheme of $\mathrm{X}$. A great deal is known about $\mathrm{F}$ for $d=3$ or $d=4$ [2, 8, 10, 13, 28], and so our focus is on the $d\geqslant 5$ case, which has received much less attention. Altman and Kleiman [1] show that $\mathrm{F}$ is an irreducible normal local complete intersection of dimension 2d42d-4, and it is known that $\mathrm{X}$ can be recovered from $\mathrm{F}$ [7].
@@
$\mathcal{H}_{\mathrm{X}/\mathbf{P}}$$\mathcal{D}_{\mathrm{X}}$$\mathcal{H}_{\mathrm{F}/\mathbf{G}}$$\mathcal{D}_{\mathrm{F}},$$\scriptstyle{\eta_{\mathcal{H}}}$$\scriptstyle{\eta}$
@@
$\mathrm{H}^{0}(\mathcal{N}_{\mathrm{X}/\mathbf{P}})$$\operatorname{Ext}^{1}(\Omega^{1}_{\mathrm{X}},\mathcal{O}_{\mathrm{X}})$$\mathrm{H}^{0}(\mathcal{N}_{\mathrm{F}/\mathbf{G}})$$\operatorname{Ext}^{1}(\Omega^{1}_{\mathrm{F}},\mathcal{O}_{\mathrm{F}})$$\scriptstyle{d\eta_{\mathcal{H}}}$$\scriptstyle{d\eta}$
of differentials. Relying on Borel-Bott-Weil computations and hypercohomology spectral sequences associated to the Koszul resolution of $\mathcal{O}_{\mathrm{F}}$, we show that the maps $\mathrm{H}^{0}(\mathcal{N}_{\mathrm{F}/\mathbf{G}})\rightarrow\operatorname{Ext}^{1}(\Omega^{1}_{\mathrm{F}},\mathcal{O}_{\mathrm{F}})$ and $d\eta_{\mathcal{H}}$ are surjective; we then observe that
@@
using a result of Charles [7] which relates the automorphism group of $\mathrm{F}$ to the one of $\mathrm{X}$. The condition $\mathrm{H}^{0}(\Theta_{\mathrm{X}})=0$, which holds for example for Lefschetz cubics, then guarantees that both $\mathcal{D}_{\mathrm{X}}$ and $\mathcal{D}_{\mathrm{F}}$ are pro-representable. Without assuming $\mathrm{H}^{0}(\Theta_{\mathrm{X}})=0$, we show that $\eta_{\mathcal{H}}$ is an isomorphism and $\eta$ is surjective.
We should discuss the relation of our functorial approach to the work of Borcea [5] and Wehler [28]. Writing $\mathrm{X}=\mathrm{Z}(f)$ for $f\in\mathrm{H}^{0}(\mathcal{O}_{\mathbf{P}}(3))$, Borcea [5] considers the deformation of $\mathrm{F}$ given by varying $f$ in $\mathrm{H}^{0}(\mathcal{O}_{\mathbf{P}}(3))$. He checks the conditions
@@
which guarantee the completeness of the deformation [5, 28], for $d\geqslant 6$. In contrast to his and other papers [1, 9] using similar methods, we explicitly compute the decomposition of the sheaves $\Lambda^{n}\mathrm{S}^{3}\mathcal{S}$ (which occur in the Koszul resolution of $\mathcal{J}_{\mathrm{F}/\mathbf{G}}$) into Schur powers. This allows us to check the conditions (1), which play an important role in our proof, for all $d\geqslant 5$, thus extending Borcea’s result to $d=5$. In Theorem˜3.1, we use the decomposition of $\Lambda^{n}\mathrm{S}^{3}\mathcal{S}$ to express the Hilbert polynomial $\chi(\mathcal{O}_{\mathrm{F}}(n))$ of $\mathrm{F}$ in terms of the Pochhammer symbol; this generalises previous results of Altman and Kleiman [1] and Libgober [18].
@@
of locally free sheaves. The Borel-Bott-Weil theorem [4, §10], which we will use frequently in this paper, computes the cohomology of sheaves of the form
@@
Theorem 2.1 (Borel-Bott-Weil).
@@
We will in particular rely on the following standard applications, where we tacitly use the canonical isomorphism $\mathrm{S}^{\vee}=\mathcal{S}\otimes\det(\mathcal{Q})$.
@@
and write $q_{\mathrm{S}}$ and $p_{\mathrm{S}}$ for the projections of $\mathbf{P}_{\mathrm{S}}\times_{\mathrm{S}}\mathrm{F}(\mathbf{P}_{\mathrm{S}}/\mathrm{S})$ to $\mathbf{P}_{\mathrm{S}}$ and $\mathrm{F}(\mathbf{P}_{\mathrm{S}}/\mathrm{S})$, respectively. By [1, Theorem 2.17], the closed subscheme $\mathrm{F}(\mathrm{X}/\mathrm{S})\subset\mathrm{F}(\mathbf{P}_{\mathrm{S}}/\mathrm{S})$ is the zero scheme of the canonical morphism
@@
of sheaves on $\mathbf{P}_{\mathrm{S}}\times_{\mathrm{S}}\mathrm{F}(\mathbf{P}_{\mathrm{S}}/\mathrm{S})$. Of course, $\mathrm{F}(\mathbf{P}_{\mathrm{S}}/\mathrm{S})=\mathbf{G}\times\mathrm{S}$, where we view $\mathbf{G}$ as $\mathrm{F}(\mathbf{P}/\operatorname{Spec}(\mathbf{C}))$; writing $\pi:\mathbf{G}\times\mathrm{S}\rightarrow\mathbf{G}$ for the projection, we have
@@
If $\mathrm{X}=\mathrm{Z}(f)$ for $f\in\mathrm{H}^{0}(\mathcal{O}_{\mathbf{P}_{\mathrm{S}}}(3))$, then applying $p_{\mathrm{S}\ast}$ to (4) induces a section $\sigma_{f}$ of
@@
such that, invoking [1, Proposition 2.3],
@@
If $\mathrm{S}=\operatorname{Spec}(\Lambda)$ is affine, we use the abbreviation $\mathrm{F}(\mathrm{X}/\Lambda)=\mathrm{F}(\mathrm{X}/\operatorname{Spec}(\Lambda))$. We first consider $\mathrm{F}(\mathrm{X})=\mathrm{F}(\mathrm{X}/\mathbf{C})$ for a cubic $\mathrm{X}=\mathrm{Z}(f)$, $f\in\mathrm{H}^{0}(\mathcal{O}_{\mathbf{P}}(3))$. This scheme is particularly well-behaved when the singular locus of $\mathrm{X}$ is finite, see Corollary 1.4 and Proposition 1.19 of [1]:
Theorem 2.2 (Altman-Kleiman).
Let $\mathrm{X}$ be a cubic with finitely many singularities. The Hilbert scheme $\mathrm{F}=\mathrm{F}(\mathrm{X})$ is of pure dimension 2d42d-4; moreover, $\mathrm{F}$ is reduced for $d\geqslant 4$.
@@
Here the sum is taken over all partitions $\lambda$ of $nm$ with at most $n$ parts, and the numbers $a^{\lambda}_{n,m}$ can be expressed in terms of generalized Kostka numbers [19, I §8]. For small $n$ and $m$, these coefficients are relatively easy to compute; we find
@@
By applying Lemma˜2.1, $\mathcal{S}^{\vee}=\mathcal{S}\otimes\det(\mathcal{Q})$, and the Pieri rule, we obtain
@@
For certain classes of complete intersections (including cubics of dimension $d\geqslant 6$), the latter result was obtained by Borcea [5, §5]; our approach is similar to his, but Borcea does not explicitly compute the plethysms of Lemma˜2.1 — by employing weight considerations, he instead proves a vanishing theorem (which, by Proposition˜2.1, does not hold for $d=5$).
@@
We recall now some well-known general facts about functors of Artin rings, and explain our notation; for us an Artin ring is a local $\mathbf{C}$-algebra which is finite over $\mathbf{C}$. For a functor of Artin rings $\mathcal{F}$, we denote by $t_{\mathcal{F}}=\mathcal{F}(\mathbf{C}[\varepsilon])$ the tangent space of $\mathcal{F}$, and if $\varphi:\mathcal{F}\rightarrow\mathcal{G}$ is a functorial morphism, we refer to
@@
as the differential of $\varphi$. For future reference, we state the following result [23, Remark 2.3.8]:
@@
The local moduli functor $\mathcal{D}_{\mathrm{S}}$ of a projective scheme $\mathrm{S}$ takes an Artin ring $\Lambda$ to the set $\mathcal{D}_{\mathrm{S}}(\Lambda)$ of isomorphism classes of deformations of $\mathrm{S}$ over $\Lambda$. It has three basic properties:
@@
(iii) If $\mathrm{S}$ is reduced, then $t_{\mathcal{D}_{\mathrm{S}}}=\operatorname{Ext}^{1}(\Omega^{1}_{\mathrm{S}},\mathcal{O}_{\mathrm{S}})$; if $\mathrm{S}$ is also a local complete intersection, then $\operatorname{Ext}^{2}(\Omega^{1}_{\mathrm{S}},\mathcal{O}_{\mathrm{S}})$ is an obstruction space for $\mathcal{D}_{\mathrm{S}}$.
We refer to Theorem 2.4.1, Proposition 2.4.8, and Corollary 2.6.4 of [23]; (i) is originally due to Schlessinger [21, Proposition 3.10], see [11, Proposition 4] for (ii). For a closed subscheme $\mathrm{Z}$ of $\mathrm{S}$, the Hilbert functor of $\mathrm{S}$ induces a functor of Artin rings $\mathcal{H}_{\mathrm{Z}/\mathrm{S}}$ (the local Hilbert functor), which takes an Artin ring $\Lambda$ to the set $\mathcal{H}_{\mathrm{Z}/\mathrm{S}}(\Lambda)$ of deformations of $\mathrm{Z}$ in $\mathrm{S}$ over $\Lambda$. By the existence of the Hilbert scheme of $\mathrm{S}$, $\mathcal{H}_{\mathrm{Z}/\mathrm{S}}$ is pro-representable and $t_{\mathcal{H}_{\mathrm{Z}/\mathrm{S}}}=\mathrm{H}^{0}(\mathcal{N}_{\mathrm{Z}/\mathrm{S}})$. It is related to $\mathcal{D}_{\mathrm{Z}}$ by a forgetful morphism
@@
In view of the description (5) of the Fano scheme as a zero scheme, we will need some specific results about deformations of zero schemes of sections.
Lemma 2.3.
Consider a local scheme $\operatorname{Spec}(\Lambda)$, a locally free sheaf $\mathcal{F}$ on $\mathrm{S}\times\operatorname{Spec}(\Lambda)$, and a section $\sigma\in\mathrm{H}^{0}(\mathcal{F})$. If $\sigma|_{\mathrm{S}}\in\mathrm{H}^{0}(\mathcal{F}|_{\mathrm{S}})$ is regular, then $\mathrm{Z}(\sigma)\rightarrow\operatorname{Spec}(\Lambda)$ is flat.
In particular, the morphism $\mathrm{Z}(\sigma)\rightarrow\operatorname{Spec}(\Lambda)$ is a deformation of $\mathrm{Z}(\sigma|_{\mathrm{S}})$ in $\mathrm{S}$ over $\Lambda$. After taking an open affine cover of $\mathrm{S}\times\operatorname{Spec}(\Lambda)$ trivializing $\mathcal{F}$, this is a consequence of the equational criterion for flatness [23, Example A.12]. For the rest of this section, we assume that $\mathrm{S}$ is smooth, and $\mathcal{F}$ is a locally free sheaf on $\mathrm{S}$.
@@
We refer to [23, Remark 3.2.10] for (i); (ii) is a consequence of the standard identification $\mathcal{H}_{\mathrm{Z}/\mathrm{S}}(\mathbf{C}[\varepsilon])=\mathrm{H}^{0}(\mathcal{N}_{\mathrm{Z}/\mathrm{S}})$ (see for instance the proof of [23, Proposition 3.2.1]), and Lemma˜2.3 for $\Lambda=\mathbf{C}[\varepsilon]$.
Finally, we consider the projection $\varphi:\mathrm{S}\times\mathrm{H}^{0}(\mathcal{F})\rightarrow\mathrm{S}$. There is a tautological section $\zeta\in\mathrm{H}^{0}(\varphi^{\ast}\mathcal{F})$ such that $\zeta(s,\sigma)=\sigma(s)$ for every point $(s,\sigma)$ of $\mathrm{S}\times\mathrm{H}^{0}(\mathcal{F})$; in particular, $\zeta|_{\mathrm{S}\times\{\sigma\}}=\sigma$. Let $\mathcal{Z}=\mathrm{Z}(\zeta)$ be its zero scheme, and
@@
be the projection. Then $\pi^{-1}(\sigma)=\mathrm{Z}(\sigma)$, and if $\sigma$ is regular, then Lemma˜2.3 implies that $\pi$ is flat in a neighbourhood of $\sigma$, thus inducing a deformation of $\mathrm{Z}=\mathrm{Z}(\sigma)$. The following result [28, Theorem 1.5] gives a criterion for the completeness of the latter deformation.
Lemma 2.5 (Wehler).
@@
Using Schubert calculus, Altman and Kleiman [1, Proposition 1.6] prove that the Plücker degree of $\mathrm{F}$ is given by
@@
In the special case $d=3$, this is a theorem of Fano [10, §2]. It is thus a natural question to determine, more generally, the Hilbert polynomial
$\chi(\mathcal{O}_{\mathrm{F}}(n))=\sum_{k=0}^{2d-4}\frac{n^{k}}{k!}\int_{\mathrm{F}}c_{1}(\mathcal{O}_{\mathrm{F}}(1))^{k}\cap\mathrm{Td}(\mathrm{F}).$
Altman and Kleiman (and, independently, Libgober [18, §2]) show that for $d=3$, we have
@@
In this section, we use Lemma˜2.1 to express the Hilbert polynomial $\chi(\mathcal{O}_{\mathrm{F}}(n))$, for any dimension $d$, in terms of the Pochhammer symbol.
@@
Using Lemma˜2.1, it suffices to describe the Hilbert polynomial of $\Sigma^{\mu_{1},\mu_{2}}\mathcal{S}$ for any $\mu_{1}\geqslant\mu_{2}$. We now establish the equality
@@
$\chi(\Sigma^{n^{d}}\mathcal{Q}\otimes\Sigma^{\mu_{1},\mu_{2}}\mathcal{S})=\dim\Sigma^{n^{d},\mu_{1},\mu_{2}}\mathrm{V}$
@@
by the Weyl dimension formula. In particular, combining (7) with Lemma˜2.1, we obtain
@@
A result of Schlömlich [22, §3] explicitly describes the coefficients of
@@
${d\brack k}=(-1)^{d-k}\sum_{m=0}^{d-k}\binom{d-1+m}{k-1}\binom{2d-k}{d+m}\sum_{n=0}^{m}\frac{(-1)^{n}n^{d-k+m}}{n!(m-n)!}.$
Combining this with Theorem˜3.1, we can express the coefficients of $\chi(\mathcal{O}_{\mathrm{F}}(n))$ as sums of binomial coefficients; this shows in particular that our expression for $\chi(\mathcal{O}_{\mathrm{F}}(n))$ is a polynomial of degree 2d42d-4, as it should be.
@@
Theorem˜3.1 and Kodaira vanishing
@@
allow one to compute the dimension of the space $\mathrm{H}^{0}(\mathcal{J}_{\mathrm{F}/\mathbf{G}}(n))$ of global sections of $\mathcal{O}_{\mathbf{G}}(n)$ vanishing on $\mathrm{F}$. Indeed, for $d\geqslant 4$, Debarre and Manivel [9, Theorem 4.1] prove that $\mathrm{H}^{1}(\mathcal{J}_{\mathrm{F}/\mathbf{G}}(n))=0$ for $n\geqslant 0$. There is thus an exact sequence of the form
$0\rightarrow\mathrm{H}^{0}(\mathcal{J}_{\mathrm{F}/\mathbf{G}}(n))\rightarrow\mathrm{H}^{0}(\mathcal{O}_{\mathbf{G}}(n))\rightarrow\mathrm{H}^{0}(\mathcal{O}_{\mathrm{F}}(n))\rightarrow 0.$
@@
Writing $\mathrm{X}_{d}$ to indicate the dimension $d$ of the cubic $\mathrm{X}$, we have
@@
In fact, any deformation $\mathfrak{X}\subset\mathbf{P}_{\Lambda}$ of $\mathrm{X}$ in $\mathbf{P}$ over an Artin ring $\Lambda$ is a cubic: there exists a section $f_{\Lambda}$ of $\mathcal{O}_{\mathbf{P}_{\Lambda}}(3)$ extending $f$, such that $\mathfrak{X}=\mathrm{Z}(f_{\Lambda})$ [29, Theorem 1].
@@
As the vanishing of $\mathrm{H}^{0}(\Theta_{\mathrm{X}})$ guarantees the pro-representability of $\mathcal{D}_{\mathrm{X}}$, we are led to study $\mathrm{H}^{0}(\Theta_{\mathrm{X}})$. It is well-known that $\mathrm{H}^{0}(\Theta_{\mathrm{X}})=0$ when $\mathrm{X}$ is smooth (see [14, §5], [16, Lemma 14.2]). We now extend this result to the class of Lefschetz cubics in the sense of [8, Definition 5.1]; apart from smooth cubics, this class consists of the simplest singular cubics: those whose singular locus consists of a single node. Observe that $\mathrm{H}^{0}(\Theta_{\mathrm{X}})$ is the kernel of the derivative
@@
$\mathrm{H}^{0}(\Theta_{\mathbf{P}}|_{\mathrm{X}})$$\mathrm{H}^{0}(\mathcal{O}_{\mathrm{X}}(3))$$\mathrm{H}^{0}(\Theta_{\mathbf{P}})$$\mathrm{H}^{0}(\mathcal{O}_{\mathbf{P}}(3))/(f)$$\scriptstyle{df}$
$\scriptstyle\thicksim$
$\scriptstyle\thicksim$
is given by $df(\sum L_{i}\partial_{i})=\sum L_{i}\partial_{i}f\mod(f)$. We can thus view $\mathrm{H}^{0}(\Theta_{\mathrm{X}})$ as the subspace of $\mathrm{H}^{0}(\Theta_{\mathbf{P}})$ consisting of all $\sum L_{i}\partial_{i}$ such that